# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

### v1

@@ IV-B1 EA-Based Methods
EAs have a long-standing history of jointly evolving robot morphology and control. In these methods, each individual in the population encodes a complete agent, encompassing both the body structure and the controller parameters. The evolutionary loop evaluates fitness and applies operators such as selection, mutation, and crossover to explore the combined design space.
Early work by Sims [sims1994evolving, sims1994competition] evolved 3D virtual creatures with embedded neural controllers, demonstrating the emergence of locomotion and showing that graph-based encodings naturally yield modular and symmetric morphologies. Subsequent research expanded this paradigm through diverse genotype representations, including developmental encodings [bongard2001repeated, brodbeck2015morphological], CPPNs [auerbach2011evolving, tanaka2022co], and learned morphological representations [zhao2025concurrent]. Recent efforts have introduced novel evolutionary search mechanisms. Morphological innovation protection [cheney2018scalable] mitigates premature convergence by preserving newly evolved traits, while other studies incorporate heuristic search [ha2018computational], biologically inspired strategies [davis2023subtract], and neuroscience-inspired approaches [ferigo2025totipotent]. Multi-objective EAs, such as NSGA-II [deb2002fast], have been utilized to balance competing design objectives and constraints. For example, minimizing energy consumption and mission time jointly in morphing aerial vehicles [bergonti2024co].
EA-based methods excel at discovering creative, unexpected morphologies due to their global and relatively unconstrained search. In soft robotics, they have been widely used to automatically generate complex geometries and behaviors [cheney2013unshackling, hiller2010evolving, cheney2014electro, hiller2012automatic, kriegman2021fractals]. However, scalability remains a major limitation: population-based evaluation is computationally expensive—often requiring thousands of simulations per generation—and convergence is typically slow, mirroring the efficiency bottlenecks observed in bi-level co-design.
@@ IV-B2 RL-Based Methods
Early methods incorporated morphology directly into the policy representation. For example, [ha2019reinforcement] used a population-based policy-gradient method (REINFORCE) to update control and morphology simultaneously. Similarly, [chen2018hardware] embedded body parameters into the computational graph of an RL policy, enabling joint optimization of physical parameters and neural network weights. [schaff2019jointly] co-optimized a differentiable distribution over morphology parameters alongside a universal controller; this framework was later extended by [schaff2022n] for more complex embodiments. Other works, such as [whitman2020modular2], employed Q-learning to guide modular manipulator design, while [luck2020data, luck2021robot] combined RL with a learned Q-value surrogate to predict the agent’s task performance.
A complementary paradigm treats design modifications themselves as actions within the RL framework. Under this formulation, the agent concurrently learns to control and perform morphological “design actions”. A representative work is Transform2Act [yuan2022transformact], which introduces a transform-and-control paradigm that formulates design optimization as a conditional policy over both control and morphology, with policy-gradient updates and experience sharing across designs for improved sample efficiency. Subsequent extensions incorporated curriculum learning [wang2023curriculum], multi-task pretraining [wang2023preco], symmetry priors [dong2023symmetry], hierarchical modeling [sun2023co], competition mechanisms [huang2024competevo], and temporal credit assignment [lu2025bodygen], with applications ranging from embodied manipulation [liu2023learning, kulz2025design] to freeform modular robot design [li2024reinforcement]. Moreover, real-world considerations have also motivated sim-to-real approaches. [schaff2022soft] showed that co-designed soft robots can effectively navigate complex terrains. MORPH [he2024morph] co-optimized a neural proxy of hardware alongside the control policy, enabling gradient-based updates even with non-differentiable physics.
@@ IV-B3 Physics-Based Methods
Another approach to single-level co-design leverages physics-based optimization. If the simulator, analytic model, or learned dynamics is differentiable, gradients can be computed with respect to both the controller and body parameters. Task losses can then be backpropagated through time using standard optimizers (Grad. Opt.), such as ADAM [adam] or L-BFGS [zhu1997algorithm]. Compared with black-box EA search or gradient approximations in RL, these approaches exhibit greater sample efficiency.
Early efforts relied on parametric analytic models [spielberg2017functional, ha2018computational] and adjoint methods [desai2018interactive] to model the complex relationships between morphology and control parameters. Subsequent research introduced stochastic programming combined with trajectory optimization [bravo2020one, bravo2024engineering], improving scalability across tasks and environments. More recent methods incorporate hardware and environmental constraints [ghansah2023humanoid, vaish2024co, sartore2022optimization]. Most analytic approaches assume fixed body topologies, primarily due to modeling challenges. Differentiable simulators address this limitation by providing end-to-end gradients through various physical processes, such as soft-body [du2021diffpd, ma2021diffaqua, qiao2021differentiable, hu2019chainqueen, hu2019difftaichi], rigid-body [freeman2021brax, xu2021end, degrave2019differentiable], and differentiable rendering [murthy2020gradsim, ma2021risp]. These processes enable gradients to propagate through more expressive design spaces. Built on this foundation, topology-agnostic parameterizations, such as deformation-based [xu2021end, cai2023task2morph, liu2023learning, lirobottool2023] and density-based schemes [yuhn20234d, matthews2023efficient, sato2023topology, sato2025computational], have been proposed for agents with complex geometries. Additional mechanisms include particle sets [spielberg2021co, matthews2023efficient], Wasserstein barycenters [ma2021diffaqua, wang2023softzoo], and CPPN-based encodings [cochevelou2023differentiable].
@@
Krvcah et al. [krvcah2010solving]
2010
Virtual
Creature
Simulation
{ODE}
EA (NEAT) ANN EA (NEAT) Nested Graph Loco.&Navi. Task+Novelty
Lehman et al. [lehman2011evolving]
2011
Virtual
Creature
Simulation
{ODE}
EA (NSGA‑II) ANN EA (NSGA‑II) Nested Graph Loco.&Navi. Task+Novelty
@@ IV-B4 Diversity-Based Methods
A persistent challenge in single-level co-design is the premature convergence of both morphology and controller. Diversity-based methods address this issue by optimizing a repertoire of distinct, high-quality body–brain pairs instead of focusing on a single global optimum. Early work [krvcah2010solving, lehman2011evolving] demonstrated this principle by employing Novelty Search [doncieux2019novelty] with local competition within a three-dimensional morphology space (height, mass, and joint count), showing that localized selection pressures help preserve morphological diversity.
The subsequent line built upon Quality-Diversity (QD) algorithms [pugh2016quality, cully2017quality]. Canonical QD methods target behavioral diversity [batra2024quality, nilsson2021policy, flageat2024enhancing, pierrot2022multi, faldor2023map, pierrot2022diversity]. For example, MAP-Elites [mouret2015illuminating] discretizes a user-defined behavior descriptor (BD) space into cells and stores the best-performing solution within each. In the context of ECD, morphology descriptors (MDs) are introduced to index MAP-Elites archives. For instance, [nordmoen2020quality, nordmoen2021map] discretized a two-dimensional MD space (total voxels and active voxels) and stored the fastest walker in each cell, achieving greater robustness and diversity compared to objective-only EAs. Recent works extend this framework to legged robots and manipulators [howard2020diversity, xie2024map, norstein2023effects], or distill the archive into a universal controller to stabilize the co-design process [mertan2025controller]. Moreover, dual-map approaches separately represent morphology and controller spaces [zardini2021seeking], whereas 3B-QD [nadizar2025enhancing] maintains three independent archives for morphology, control, and behavior, further enhancing the illumination of the agent co-design space.
Diversity-based methods accelerate convergence by avoiding repeated controller optimization for each new morphology. Remaining challenges include the automatic discovery of informative descriptors, scaling to high-resolution material-property spaces [pigozzi2023factors], and providing theoretical guaranties for niche-optimal solutions. Comparative studies of QD approaches in ECD are available in [mkhatshwa2023impact, mkhatshwa2025body].
@@ IV-C Generative Co-Design
Generative co-design leverages automated generative models to capture the relationship among morphology, task, and performance (evaluated by learning a controller), thereby enabling the optimization of morphology within latent or rule-defined spaces, as illustrated in Fig. 4 (c). We further categorize the generative co-design framework into three topics: 1) Rule-Based Generation, 2) Latent-Based Generation, and 3) Large-Model-Based Generation. Each class employs a distinct generative representation and search strategy to explore morphological design spaces, as shown in Table V.
@@ IV-C1 Rule-Based Generation
Early work along this line typically utilized graph-based generative encodings [sims1994competition, sims1994evolving, pilat2008creature] and L-systems [hornby2001evolution, hornby2001evolving, hornby2003generative] to jointly generate morphology and controller parameters, which were then evolved to create task-oriented agents. More recent approaches combine grammars with advanced search strategies. Robogrammar [zhao2020robogrammar] and its extension [zhao2025multi], for example, encode robots as graphs, enforce valid assemblies through a graph grammar, and couple heuristic search with a model-based controller to efficiently identify high-performing designs. Zhao et al. [zhao2022automatic] extended this framework to aerial vehicles by combining discrete grammar exploration with a differentiable simulator, enabling both rediscovery of canonical drones as well as the discovery of novel configurations. Zharkov et al. [zharkov2024synergizing] further applied grammar-based linkage graphs to co-design underactuated, tendon-driven grippers.
@@
Rules effectively reduce the search space to plausible designs, explicitly encode prior knowledge such as symmetry and connectivity, and enhance manufacturability and modularity. While they are flexible and interpretable, these rules depend on expert-crafted grammars, which can introduce bias into the search process and limit the exploration of novel or fine-grained morphological variations.
IV-C2 Latent-Based Generation
Variational Autoencoders (VAEs) learn a continuous latent representation of morphology and, in some cases, control parameters, facilitating sample-efficient optimization. LABO [pan2021emergent] embeds high-dimensional hand morphologies and controller parameters into a VAE latent space and performs BO within that space, thereby reducing dimensionality and sample complexity for grasping tasks. GLSO [hu2023glso] extends this idea to graph-structured robots by training a graph-VAE on grammar-generated topologies and optimizing designs using BO. Liu et al. [liu2025terrain] proposed a Grammar-VAE that maps discrete morphologies into a continuous latent space and integrated differential evolution with a neural decoder, allowing the generative model to be reused across tasks. MorphVAE [song2024morphvae] further conditions latent sampling on task embeddings and introduces a “continuous natural selection” strategy to improve diversity and performance.
Diffusion Models have recently emerged as powerful generative priors for embodied co-design. DiffuseBot [wang2023diffusebot] integrates differentiable physics into the denoising process, using simulation feedback at each step to bias generation toward high-performing designs. Dynamics-Guided Diffusion Models (DGDM) [xu2024dynamics] replace task-specific training data with physics-based guidance, enabling generalization to new manipulation tasks and novel geometries.
GFlowNets provide an alternative generative formulation in which morphology–controller pairs are sampled with probability proportional to their simulated performance. Recent works by [nagiredla2024robonet, nagiredlate, nagiredlacogent] demonstrated that GFlowNets can produce diverse, high-quality designs while learning a structured generative policy over complete embodied agents.
Latent generative models effectively capture complex, free-form geometries, exploit correlations among design variables, and typically offer significant gains in sample efficiency when compared to direct search methods. Once trained, a latent model can be reused across various tasks or paired with different controllers. However, training generative models typically requires large datasets and simulation efforts. Moreover, latent spaces may decode to invalid or unrealistic morphologies without appropriate constraints. Achieving the right balance between fidelity and diversity remains challenging.
IV-C3 Large-Model-Based Generation
Large pretrained models—including Large Language Models (LLMs) [naveed2025comprehensive] and Vision–Language Models (VLMs) [zhang2024vision]—have recently been introduced into ECD as generative priors or intelligent search operators. These models leverage broad knowledge and powerful generative capabilities to propose, mutate, or refine robot morphologies and control strategies.
A growing line of work uses LLMs as intelligent mutation operators or search heuristics. OpenELM [lehman2023evolution] demonstrates that a LLM trained on code can act as a high-quality mutation operator in genetic programming. In Sodarace experiments, the LLM repeatedly mutated code that generates robot walkers, producing thousands of novel and functional agents without prior robotics data; the resulting dataset was then used to train a conditional generative model that maps terrain descriptions to robot code. Building on this paradigm, LASeR [song2025laser] integrates a LLM into the optimization loop via a reflection mechanism that summarizes past search histories and prompts the LLM to generate improved design variants. RoboMoRe [fang2025robomore] further investigates LLM-driven co-design by asking the LLM to jointly propose candidate morphology–reward pairs during a coarse search, after which promising candidates undergo alternating refinement. RoboMorph [qiu2024robomorph] represents each robot with a structured grammar and relies on LLM reasoning to navigate vast combinatorial spaces, iteratively improving morphologies through feedback.
LLMs have also been utilized for human–AI collaborative design. Stella et al. [stella2023can] have shown that a LLM can provide high-level conceptual guidance and partial specifications for a tomato-harvesting gripper, illustrating its potential in early-stage design. Beyond text-based reasoning, vision models are beginning to appear in ECD. VLM-Engineer [gao2025vlmgineer] combines VLMs with evolutionary search to jointly design robotic tools and associated control policies. This line of work suggests future workflows in which a VLM first generates sketches or 2D concept designs from a task description, which are then lifted into full 3D morphologies.
Large models offer several advantages for ECD: they encapsulate extensive engineering and commonsense knowledge, generate creative or unconventional designs, support semantic prompting (e.g., “design a robot capable of climbing a ladder”), and diversify the search space. Their in-context learning abilities make them powerful assistants for bridging high-level human intent with concrete design proposals [makatura2023can, ma2024llm]. However, outputs from LLMs and VLMs are not guaranteed to be feasible, optimal, or physically consistent. Their behavior is often sensitive to prompt phrasing, and unguided LLM-driven searches can be highly sample-inefficient. Vision-based 3D generation is still in its infancy, and the reliance on proprietary foundation models raises additional concerns regarding reproducibility, transparency, and computational costs.
@@ IV-C3 Large-Model-Based Generation
IV-D Open-Ended Co-Design
Open-ended co-design aims for the continual or unbounded improvement of agents by enabling their bodies, controllers, and environments to co-evolve. These systems promote continuous innovation by enabling agents and environments to mutually generate new challenges and adaptations, as shown in Fig. 4 (d). We categorize open-ended co-design into two primary subtopics: 1) Brain–Body–Environment Co-Evolution and 2) Developmental Evolution, as summarized in Table VI.
IV-D1 Brain-Body-Environment Co-Evolution
This approach treats the task environment as a co-optimized component alongside morphology and control. By evolving environments (or task curricula) together with agents, it creates continuously shifting challenges that prevent premature convergence and promote both behavioral and morphological improvement. The environment thus functions simultaneously as a generator of new selective pressures and as a filter for successful morphology–controller adaptations.
Foundational work in this direction is the Paired Open-Ended Trailblazer (POET) framework [wang2019poet, wang2020enhanced]. POET maintains a population of environment–controller pairs, simultaneously mutating environments and evolving agents. Policies that succeed in one environment may be transferred to other contexts. Stensby et al. [stensby2021co] extended POET to jointly evolve morphologies and controllers in tandem with their environments. Their results indicate that learned environment progressions yield greater morphological diversity than handcrafted curricula and reduce premature convergence [eguiarte2024premature].
Recent work has introduced RL-driven co-evolution of the brain, body, and environment. Ao et al. proposed MECE [ao2023curriculum], which formulates a curriculum in which both morphology and the environment evolve according to learned RL policies. This produces progressively challenging and structured tasks that drive sustained adaptation. Another emerging trend augments environment evolution with generative models. For example, LLM-POET [aki2024llm] replaces manual environment mutation with a LLM that generates increasingly complex environments for a population of soft-robot agents. As agents improve, the LLM proposes novel tasks that escalate difficulty, realizing an open-ended loop of continual challenge and innovation.
IV-D2 Developmental Evolution
Developmental co-evolution draws inspiration from evolutionary developmental biology (Evo-Devo) and developmental robotics, where genotypes encode growth processes that shape both morphology and control over an agent’s lifetime (ontogeny) [asada2001cognitive, bongard2003evolving, lungarella2003developmental, kriegman2018morphological, van2021comparing, ivanchenko2003developmental]. Open-endedness emerges from developmental plasticity [higashinaka2024evolution]: agents are evaluated across multiple developmental stages or allowed to modify their bodies during their lifetime, while evolution optimizes the developmental rules themselves.
Early research has explored developmental encodings, including L-systems [rieffel2014growing], grammar rules, and neural networks [pontes2022single, joachimczak2016artificial]. Bongard [bongard2003evolving] demonstrated that artificial ontogeny can generate hierarchical, repeated structures and controllers that emerge as bodies grow, yielding more structured solutions than direct genotype-phenotype encodings. The other research line focuses on morphological development. Here, developmental rules adjust parameters such as leg angles or actuator amplitudes during learning [bongard2011morphological, bongard2010utility, naya2020experiment], or progressively “unfreeze” degrees of freedom to facilitate skill acquisition [lungarella2002adaptivity, berthouze2004motor, savastano2012incremental]. Later work introduced mechanisms for growing new body parts during an agent’s lifetime [arita2016alife, benureau2022morphological, jaafar2025eco], providing additional adaptive flexibility. For real-world deployment, virtual growth is typically halted once the simulated agent reaches the physical dimensions of the hardware robot, after which the learned controller is transferred to the physical platform [vujovic2017evolutionary, brodbeck2015morphological]. For comprehensive overviews of developmental evolution, we refer readers to [naya2021morphological, wang2021survey]. Overall, open-ended co-design is not yet a unified methodology but rather a collection of emerging ideas that leverage ontogeny, plasticity, and staged learning, and it awaits further exploration.

### v2

@@ IV-B1 EA-Based Methods
EAs have a long-standing history of jointly evolving robot morphology and control. In these methods, each individual in the population encodes either a complete agent or only its morphology, along with a simplified controller. The evolutionary loop evaluates fitness and applies operators such as selection, mutation, and crossover to explore the combined design space.
Early work by Sims [sims1994evolving, sims1994competition] evolved 3D virtual creatures with embedded neural controllers, demonstrating the emergence of locomotion and showing that graph-based encodings naturally yield modular and symmetric morphologies. Subsequent research expanded this paradigm through diverse genotype representations, including developmental encodings [bongard2001repeated, brodbeck2015morphological], CPPNs [auerbach2011evolving, tanaka2022co], and learned morphological representations [zhao2025concurrent]. Recent efforts have introduced novel evolutionary search mechanisms. Morphological innovation protection [cheney2018scalable] mitigates premature convergence by preserving newly evolved traits, while other studies incorporate heuristic search [ha2018computational], biologically inspired strategies [davis2023subtract], and neuroscience-inspired approaches [ferigo2025totipotent]. Moreover, multi-objective EAs, such as NSGA-II [deb2002fast], have been utilized to balance competing design objectives and constraints. For example, minimizing energy consumption and mission time jointly in morphing aerial vehicles [bergonti2024co].
EA-based methods excel at discovering novel morphologies due to their global and relatively unconstrained search capabilities. These methods have been widely used to generate complex geometries and behaviors [cheney2013unshackling, hiller2010evolving, cheney2014electro, hiller2012automatic, kriegman2021fractals, sathuluri2023robust]. However, scalability remains a major limitation: population-based evaluation is computationally expensive—often requiring thousands of simulations per generation—and convergence is typically slow, mirroring the efficiency bottlenecks observed in bi-level co-design.
@@ IV-B2 RL-Based Methods
Early methods incorporated morphology directly into the policy representation. For example, [ha2019reinforcement] used a population-based policy-gradient method (REINFORCE) to update both control and morphology simultaneously. Similarly, [Chen2020HardwareAP] embedded body parameters into the computational graph of a RL policy. [schaff2019jointly] co-optimized a differentiable distribution over morphology parameters alongside a universal controller; this framework was later extended by [schaff2022n] for more complex embodiments. Other works, such as [whitman2020modular2], employed Q-learning to guide modular manipulator design.
A complementary paradigm treats design modifications as actions within the RL framework. Under this formulation, the agent concurrently learns to control and perform morphological “design actions”. A representative work is Transform2Act [yuan2022transformact], which introduces a transform-and-control paradigm that formulates design optimization as a conditional policy over both control and morphology. This approach incorporates policy-gradient updates and experience sharing across designs to improve sample efficiency. Subsequent extensions have incorporated curriculum learning [wang2023curriculum], multi-task pretraining [wang2023preco], symmetry priors [dong2023symmetry], hierarchical modeling [sun2023co], competition mechanisms [huang2024competevo], and temporal credit assignment [lu2025bodygen], with applications ranging from embodied manipulation [liu2023learning, kulz2025design] to freeform modular robot design [li2024reinforcement]. Moreover, real-world considerations have motivated sim-to-real approaches. [schaff2022soft] showed that co-designed soft robots can effectively navigate complex terrains. MORPH [he2024morph] co-optimized a neural proxy of hardware alongside the control policy, enabling gradient-based updates even with non-differentiable physics.
@@ IV-B3 Physics-Based Methods
Another approach to single-level co-design leverages physics-based optimization. If the simulator, analytic model, or learned dynamics is differentiable, gradients can be computed with respect to both the controller and body parameters. Task losses can then be backpropagated through time using gradient-based optimizers (Grad. Opt.), such as ADAM [adam] or L-BFGS [zhu1997algorithm]. Compared with black-box EA search or gradient approximations in RL, these approaches exhibit greater sample efficiency.
Early efforts relied on parametric analytic models [spielberg2017functional, ha2018computational] and adjoint methods [desai2018interactive] to model the complex relationships between morphology and control parameters. Subsequent research introduced stochastic programming in combination with trajectory optimization [bravo2020one, bravo2024engineering], thereby improving scalability across tasks and environments. More recent methods incorporate hardware and environmental constraints [ghansah2023humanoid, vaish2024co, sartore2022optimization]. Most analytic approaches assume fixed body topologies, primarily due to modeling challenges. Differentiable simulators address this limitation by providing end-to-end gradients through various physical processes, such as soft-body [du2021diffpd, ma2021diffaqua, qiao2021differentiable, hu2019chainqueen, hu2019difftaichi], rigid-body [freeman2021brax, xu2021end, degrave2019differentiable], and differentiable rendering [murthy2020gradsim, ma2021risp]. These processes enable gradients to propagate through more expressive design spaces. Built on this foundation, deformation-based [xu2021end, cai2023task2morph, liu2023learning, lirobottool2023] and density-based schemes [yuhn20234d, matthews2023efficient, sato2023topology, sato2025computational] have been proposed for agents with complex geometries. Additional mechanisms include particle sets [spielberg2021co, matthews2023efficient], Wasserstein barycenters [ma2021diffaqua, wang2023softzoo], and CPPN-based encodings [cochevelou2023differentiable].
@@ IV-B4 Diversity-Based Methods
A persistent challenge in single-level ECD is the premature convergence of both morphology and controller. Diversity-based methods address this issue by optimizing a repertoire of high-quality body–brain pairs, rather than concentrating on a single global optimum.
Early work by [krvcah2010solving, lehman2011evolving] demonstrated this principle by employing Novelty Search [doncieux2019novelty] with local competition in a three-dimensional morphology space (height, mass, and joint count), showing that localized selection pressures help preserve morphological diversity. The subsequent work built upon Quality-Diversity (QD) algorithms [pugh2016quality, cully2017quality]. Canonical QD methods target only behavioral diversity. For example, MAP-Elites [mouret2015illuminating] discretizes a user-defined behavior descriptor (BD) space into cells and stores the best-performing solution within each cell. In the context of ECD, morphology descriptors (MDs) are introduced to index MAP-Elites archives. For instance, [nordmoen2020quality, nordmoen2021map] discretized a two-dimensional MD space (total voxels and active voxels) and stored the fastest walker in each cell, achieving greater robustness and diversity compared to objective-only EAs. Recent works extend this framework to legged robots and manipulators [howard2020diversity, xie2024map, norstein2023effects], or distill the archive into a universal controller to stabilize the co-design process [mertan2025controller]. Moreover, dual-map approaches separately represent morphology and controller spaces [zardini2021seeking], while 3B-QD [nadizar2025enhancing] maintains three independent archives for morphology, control, and behavior, further enhancing the illumination of the agent co-design space.
Despite their effectiveness, diversity-based ECD faces several challenges, including the automatic discovery of informative descriptors, scalability to high-dimensional material and property spaces [pigozzi2023factors, mertan2024no], and the lack of theoretical guarantees for niche-optimal solutions. Comparative evaluations of QD-based approaches in ECD can be found in [mkhatshwa2023impact, mkhatshwa2025body].
@@ IV-C Generative Co-Design
Generative co-design leverages automated generative models to capture the relationships among morphology, tasks, and performance (evaluated by learning a controller). This approach enables optimization within latent or rule-defined spaces, as illustrated in Fig. 3 (c). We further categorize the generative co-design framework into three sub-topics: 1) Rule-Based Generation, 2) Latent-Based Generation, and 3) Large-Model-Based Generation. Each class employs a distinct generative representation and search strategy to explore morphology spaces, as summarized in Table V.
@@ IV-C1 Rule-Based Generation
Early work along this line typically utilized graph-based generative encodings [sims1994competition, sims1994evolving, pilat2008creature] and L-systems [hornby2001evolution, hornby2001evolving, hornby2003generative] to jointly generate morphology and controller parameters, which were then evolved to create task-oriented agents. More recent approaches combine grammars with advanced search strategies. Robogrammar [zhao2020robogrammar] and its extension [xu2021multi] encode robots as graphs, enforce valid assemblies through a graph grammar, and couple heuristic search with a model-based controller to efficiently identify high-performing designs. Zhao et al. [zhao2022automatic] extended this framework to aerial vehicles by combining discrete grammar exploration with a differentiable simulator, enabling both rediscovery of canonical drones as well as the discovery of novel configurations. Zharkov et al. [zharkov2024synergizing] further applied grammar-based linkage graphs to co-design underactuated, tendon-driven grippers.
@@
Rule-based approaches effectively constrain the search space to plausible designs, explicitly encode prior knowledge (e.g., symmetry and connectivity), and improve manufacturability. However, their reliance on expert-crafted grammars can introduce bias and restrict the exploration of novel or fine-grained morphological variations.
@@
IV-C2 Latent-Based Generation
Variational Autoencoders (VAEs) learn continuous latent representations of morphology and, in some cases, control parameters, facilitating sample-efficient optimization. For example, LABO [pan2021emergent] embeds high-dimensional hand morphologies and controller parameters into a VAE latent space and performs BO within that space, thereby reducing dimensionality and sample complexity for grasping tasks. GLSO [hu2023glso] extends this concept to graph-structured robots by training a graph-VAE on grammar-generated topologies and optimizing designs using BO. Liu et al. [liu2025terrain] proposed a Grammar-VAE and integrated differential evolution with a neural decoder, allowing the generative model to be reused across tasks. MorphVAE [song2024morphvae] further conditions latent sampling on task embeddings and introduces a continuous natural selection strategy to improve diversity and performance.
Diffusion Models have recently emerged as powerful generative priors for embodied co-design. DiffuseBot [wang2023diffusebot] integrates differentiable physics into the denoising process, using simulation feedback at each step to bias generation toward high-performing designs. Another recent work is Dynamics-Guided Diffusion Models (DGDM) [xu2024dynamics], which replaces task-specific training data with physics-based guidance, enabling generalization to new manipulation tasks and novel geometries.
GFlowNets offer an alternative generative formulation in which morphology–controller pairs are sampled with a probability proportional to their simulated performance. Recent works by [nagiredla2024robonet], [nagiredlate], and [nagiredlacogent] have demonstrated that GFlowNets can generate diverse, high-quality designs while learning a structured policy over complete embodied agents.
Generative models can capture the inter-correlations among design variables. Once trained, a latent model can be reused across various tasks or paired with different controllers. However, training generative models typically requires large datasets and substantial simulation efforts. Moreover, latent spaces may decode to invalid or unrealistic morphologies without appropriate constraints. Achieving the right balance between fidelity and diversity remains a challenge.
IV-C3 Large-Model-Based Generation
Large pretrained models—including Large Language Models (LLMs) [naveed2025comprehensive] and Vision–Language Models (VLMs) [zhang2024vision]—have recently been introduced into ECD, by leveraging broad knowledge and powerful generative capabilities to propose, mutate, or refine morphologies and control strategies.
A growing body of work utilizes LLMs as intelligent mutation operators or search heuristics. A pioneering work is OpenELM [lehman2023evolution], which demonstrates that a LLM trained on code can act as a high-quality mutation operator in genetic programming. In the Sodarace experiments, the LLM consistently mutated code to generate robot walkers, and the resulting dataset was then used to train a conditional generative model that maps terrain descriptions to robot code. Building on this paradigm, LASeR [song2025laser] integrates a LLM into the optimization loop via a reflection mechanism that summarizes past search histories, prompting the LLM to generate improved design variants. RoboMoRe [fang2025robomore] further investigates LLM-driven co-design by instructing the LLM to propose candidate morphology–reward pairs jointly during a coarse search, after which promising candidates undergo alternating refinement. RoboMorph [qiu2024robomorph] represents each robot with a structured grammar and relies on LLM reasoning to iteratively improve morphologies through feedback.
LLMs have also been utilized for human–AI collaborative design. Stella et al. [stella2023can] have shown that a LLM can provide high-level conceptual guidance and partial specifications for a tomato-harvesting gripper, illustrating its potential in early-stage design. Beyond text-based reasoning, vision models are beginning to emerge in ECD. VLM-Engineer [gao2025vlmgineer] combines VLMs with evolutionary search to jointly design robotic tools and associated control policies. This line of work suggests future workflows in which a VLM first generates sketches or 2D concept designs from a task description, which are then translated into complete 3D morphologies.
Large models offer several advantages for ECD: they encapsulate extensive engineering and commonsense knowledge, enable creative and unconventional designs, support semantic prompting (e.g., “design a robot capable of climbing a ladder”), and diversify the search space. Their in-context learning capabilities make them effective interfaces between high-level human intent and concrete design proposals [makatura2023can, ma2024llm]. However, the outputs of LLMs and VLMs are not guaranteed to be feasible, optimal, or physically consistent, highlighting the need for stronger grounding and validation mechanisms.